Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis Mackinaw Inc. processes a base chemical...

  1. Direct Materials, Direct Labor, and Factory Overhead Cost Variance Analysis

    Mackinaw Inc. processes a base chemical into plastic. Standard costs and actual costs for direct materials, direct labor, and factory overhead incurred for the manufacture of 78,000 units of product were as follows:

    Standard Costs Actual Costs
    Direct materials 226,200 lbs. at $5.60 223,900 lbs. at $5.50
    Direct labor 19,500 hrs. at $18.20 19,950 hrs. at $18.50
    Factory overhead Rates per direct labor hr.,
    based on 100% of normal
    capacity of 20,350 direct
    labor hrs.:
    Variable cost, $4.80 $92,660 variable cost
    Fixed cost, $7.60 $154,660 fixed cost

    Each unit requires 0.25 hour of direct labor.

    Required:

    a. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Material Price Variance $
    Direct Materials Quantity Variance $
    Total Direct Materials Cost Variance $

    b. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Direct Labor Rate Variance $
    Direct Labor Time Variance $
    Total Direct Labor Cost Variance $

    c. Determine variable factory overhead controllable variance, the fixed factory overhead volume variance, and total factory overhead cost variance. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    Variable factory overhead controllable variance $
    Fixed factory overhead volume variance $
    Total factory overhead cost variance $
Answer #1

Solution a:

Direct Material Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AQ * AP = AQ * SP = SQ * SP =
223900 $5.50 $1,231,450.00 223900 $5.60 $1,253,840.00 226200 $5.60 $1,266,720.00
-$22,390.00 F -$12,880.00 F
Direct Material Price Variance Direct Material Qty variance
Direct material price variance -$22,390.00 F
Direct material quantity variance -$12,880.00 F
Direct material cost variance -$35,270.00 F

Solution b:

Direct Labor Cost Variance
Actual Cost Standard cost for actual quantity Standard Cost
AH * AR = AH * SR = SH * SR =
19950 $18.50 $369,075.00 19950 $18.20 $363,090.00 19500 $18.20 $354,900.00
$5,985.00 U $8,190.00 U
Direct Labor rate Variance Direct Labor Efficiency Variance
Direct Labor Rate variance $5,985.00 U
Direct Labor Efficiency variance $8,190.00 U
Direct labor cost variance $14,175.00 U

Solution c:


Variable factory overhead controllable variance = Actual cost of variable overhead - Standard cost of variable overhead

= $92,660 - (19500*$4.80) = - $940 F

Fixed factory overhead volume variance = Budgeted fixed overhead - Fixed overhead applied

= $154,660 - (19500 * $7.60) = $6,460 U

Total factory overhead cost variance = - $940 F + $6,460 U = $5,520 U